OpenStudy (anonymous):

I need to use Simpson's Rule to estimate the distance travelled by a cyclist. I've been given the instantaneous speed (v) in mph and the time in minutes for 10 one minute intervals. I don't know what my function should be to calculate the distance. Any help? Thanks!

OpenStudy (neer2890):

\[\int\limits_{a}^{b} ydx= \frac{ h }{ 3 }[y0+4(y1+y3+y5+....)+2(y2+y4+y6+....)+yn]\]

OpenStudy (neer2890):

where a to b is the interval h is the difference between two intervals y is the function itself and yo to yn are the values of function at each interval for e.g., if y=1+x then y0= the value of y at x=0 which is 1 so y0=1

OpenStudy (neer2890):

so here, h=1 because of 1 minute interval and function y=distance =speed/time y0 =value of y when x=0( x can be the variable you can adjust in speed/time) y1= value of y at x=1 and so on

OpenStudy (anonymous):

Thank you! Yeah, I did have all that except the y. Did you explain how I actually figure out what my function is? I don't know the function for finding distance with given instantaneous speed.

OpenStudy (neer2890):

distance =speed *time you got instantaneous speed, you got time, your function is complete.

OpenStudy (neer2890):

sorry., in above post i have written distance=speed/ time by mistaken

OpenStudy (anonymous):

Thank you!!

OpenStudy (neer2890):

you are welcome : )

OpenStudy (anonymous):

@neer2809 One more question.: Do I need to change my data for my speed to miles per minute? My time is given in minutes and my speed in miles per hour. :

OpenStudy (dan815):

yes you do

OpenStudy (anonymous):

Great!! Thank you!!

OpenStudy (dan815):

do you know where simpsons rule is coming from

OpenStudy (dan815):

i remember doing it for my numerical methods class but its all kind hazy, i remember some other approximations to integration.. rhomberg or something simponds 1/3 and 3/8 stuff like that

OpenStudy (anonymous):

It comes from using parabolas instead of straight line segments to approximate a curve.